if you were 12 years old and your brother was 19 years old how old are both of you in 48 years?
Found 5 solutions by ichudov, shakimblee, 2828344, neetu, vikash kumar verma:
Answer by ichudov(507)

To figure out your age 48 years later, you have to add 48 to your current age. So, you would be 12+48=60 years old, and your bro would be 19+48=67 years old. You would probably both be grandfathers at that age. If you want to live to old age, do not smoke and do not become fat.

Answer by shakimblee(1)
Answer by 2828344(3)

they would be 80

Answer by neetu(2)

60 and 67 years old.

Answer by vikash kumar verma(2)

It has been asked abut in future.
Your current age = 12 years
And your brother's age = 19 years
After 48 years
Your age will be 12+48=60 years
And your brother's age will be 19+48=67 years